# Weighted Aggregation of Conformity Scores for Classification

### v1

@@ 3.3 Consistency of DLCP
In the setting of DLCP, $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We will denote this by $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ respectively. In this case, the test set and $\widehat{\bm{w}}$ become dependent. We will present the result under a high probability event about the prediction set size.
@@
Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 with the setting of DLCP. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
@@ 3.4 Consistency of DLCP+
Similar as the DLCP setting, $\widehat{\bm{w}}$ depends on the test set, so the result of prediction set size is presented in a similar way.
@@
Let $\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x,\widehat{\bm{w}}))$ be the output of Algorithm 2 with the setting of DLCP+. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
@@
We note that in the setting of DLCP+, $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. We keep use $\mathcal{I}_{2}$ in the theorem to avoid long notation.
@@ 4 Experiments
We conducted an experiment to compare the performance of various score functions and their weighted combinations for prediction set evaluation. We focused on three powerful score functions: THR (Sadinle et al., 2019), APS (Romano et al., 2020), and RANK (Luo & Zhou, 2024). Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}$ was available, and we focused on the split of the dataset unseen during the training of the pretrained classifier [^1] The unseen dataset was further divided into $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$. The naming for $\mathcal{I}_{\textrm{train}}$ is not to be confused with the dataset for training. .
The three basic score functions used in our experiment were:
Least Ambiguous Set Values Classifier (THR) (Sadinle et al., 2019).
The score function of THR is defined as:
$s_{\text{THR}}(x,y)=\widehat{p}_{y}(x).$
Adaptive Prediction Set (APS) (Romano et al., 2020).
The score function of APS is defined as:
Rank-based Score Function (RANK) (Luo & Zhou, 2024).
The score function of RANK is defined as:
@@
Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$. Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size among all the compared score functions.
Experiments were conducted on the CIFAR-100 dataset (Alex, 2009) using a pretrained RepVGG model (Ding et al., 2021). Our experiments focus on this dataset because of the following facts we found in our experiments.
  1. 1.
    This dataset has a well pretrained model, RepVGG (Ding et al., 2021) as our benchmark model.
  2. 2.
    This is one of the most challenging classification tasks in computer vision because of its large number of classes.
  3. 3.
    There is no existing score functions proposed in earlier work that obviously outperforms all others.
Given these three aspects, it will be easier for us to present the effectiveness of the weighted score.
@@
The 10000 testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets. The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05 and two different ratios of $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$, namely 80:20 and 99:1. For each combination of $\alpha$ and data split ratio, 100 runs with different index splits were conducted to ensure robustness.
The main focus of the experiments was to compare the performance of the proposed weighted score function with three constituent base score functions: APS, THR, and RANK. Additionally, the proposed method was compared with two competitive baseline score functions, RAPS (Angelopoulos et al., 2021) and SAPS (Huang et al., 2024).
Figure 2 displays the boxplot of the coverage and size of our proposed method alongside five other score functions at a significance level of $\alpha=0.01$. To further analyze the performance across different significance levels, Figure 3 presents a plot of the size against coverage for the various score functions, with $\alpha$ values ranging from 0.01 to 0.05. To ensure a fair comparison and guarantee the coverage property, the first data split method, VFCP, was employed for all methods. Our method achieves the guaranteed coverage with the smallest prediction set size among all compared score functions, demonstrating its ability to optimize efficiency while maintaining the desired coverage level.
@@
Figure 3: Comparison of size vs. coverage for different score functions as well as our proposed method across $\alpha$ values ranging from 0.01 to 0.05. The results indicate that our weighted combination method (shown in red) consistently outperforms the three constituent basic score functions (APS, THR, RANK) as well as other competitive baseline score functions (RAPS, SAPS) in terms of achieving coverage with smaller prediction set sizes.
To investigate the impact of various data split methods (Section 2.4), we further compare the performance of our weighted score function under different split approaches. Additionally, we include the results of the five individual score functions using the VFCP split method. Table 1 offers a comprehensive overview of the coverage and size results at significance levels of $\alpha=0.01$ and 0.05. These findings demonstrate the effectiveness of the different split strategies.
@@
Weighted (VFCP) 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
Weighted (EFCP) 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
Weighted (DLCP) 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
Weighted (DLCP+) 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
@@
Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset. Results are shown as mean (standard deviation).
@@ 5 Related Work
We would like to underscore the connections between our work and existing studies on conformal prediction, particularly those that delve into model aggregation and calibration. Regarding model aggregation, Yang & Kuchibhotla (2024) introduce two selection algorithms aimed at minimizing the width of the prediction interval by aggregating and selecting from regression estimators. Stutz et al. (2023) further extends conformal prediction to scenarios where the ground truth label is ambiguous, employing predictions from multiple labels to construct valid prediction sets.
In the context of model calibration, temperature scaling (Guo et al., 2017) is utilized as a post hoc method to adjust classifier confidence levels. Additionally, Rahimi et al. (2020) endeavors to calibrate predicted probabilities from deep learning models while maintaining the rank order of these probabilities. Dabah & Tirer (2024); Xi et al. (2024) empirically demonstrate that applying post hoc calibration prior to conformal prediction can reduce model overconfidence but tends to increase the size of the prediction interval. Lin et al. (2021) addresses calibration in regression settings, leveraging kernel regression and localized conformal prediction based on a pretrained neural network mean estimator to create valid and discriminative prediction intervals.
However, the original split conformal method (Papadopoulos et al., 2002) only utilizes a portion of the data for prediction, leading to reduced data efficiency. To address this limitation, several approaches have been proposed to enhance data utilization while maintaining computational efficiency.
Aggregation methods have emerged as a popular approach to improve data utilization as well as prediction efficiency in conformal prediction. Carlsson et al. (2014) extended inductive conformal prediction (ICP) by aggregating predictions from multiple ICPs, reducing variance and improving efficiency, albeit more conservatively than ICP. The cross-conformal prediction Vovk (2015); Vovk et al. (2018) presents a hybrid of inductive conformal prediction and cross-validation, offers computational efficiency comparable to ICP while producing valid predictions empirically. The validity of aggregated conformal predictors is not fully understood and requires additional assumptions; a revised definition is proposed Linusson et al. (2017) to achieve approximate validity conditional on the underlying learning algorithm’s properties. Johansson et al. (2014) demonstrated that random forests as the underlying model for conformal regression outperform others in terms of efficiency, using out-of-bag estimates instead of a separate calibration set. BARBER et al. (2021) introduced the jackknife+ method, which employs leave-one-out cross-validation to compute nonconformity scores for each example in the calibration set, enabling the use of all available data for both training and calibration. Kim et al. (2020) proposed a similar approach using K-fold cross-validation, where the model is trained on multiple folds of the data, and the predictions from each fold are combined to create a valid prediction set. These methods strike a balance between data efficiency and computational complexity by reusing data for both training and calibration. Recently, conformal prediction has been extended to handle distribution shifts Tibshirani et al. (2019); Gibbs et al. (2023); Prinster et al. (2024), relaxing the assumption of exchangeability. This advancement allows for the application of conformal prediction to a wider range of real-world scenarios, such as data with dependencies Chernozhukov et al. (2021); Luo et al. (2023); Nettasinghe et al. (2023); Luo & Colombo (2024).
It is worth noting that there is an inherent connection between model aggregation and our approach of weighting non-conformity scores, even though our method does not directly address the scenario of combining multiple models to find an optimal aggregation.
To illustrate this connection, consider a scenario with three classifiers, including an oracle classifier making perfect predictions. A model aggregation method such as boosting (Mukherjee & Schapire, 2010) would assign a weight of 1 to the oracle classifier, disregarding the others. Similarly, when weighting non-conformity scores, even without a perfect classifier, the resulting conformal prediction sets will be equivalent to applying the THR score to the oracle classifier. The optimally chosen weighted non-conformity score can identify and prioritize the most informative and reliable scores, mimicking an oracle classifier. This equivalence highlights our method’s potential to achieve optimal performance by leveraging multiple non-conformity scores, even without a perfect classifier, approximating the performance of an ideal classifier through appropriate weight assignment.
6 Conclusion
In conclusion, our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedures. This adaptability allows our method to adjust to the specific characteristics of the dataset, potentially outperforming methods that rely on a single score function. Our theoretical analysis establishes finite-sample guarantees for coverage, ensures the validity of the prediction sets, and derives oracle inequalities that quantify the expected set size of our approach relative to an optimal weight.
Experiments on the CIFAR-100 dataset demonstrate the superior performance of our weighted approach compared to single-score methods, consistently achieving the guaranteed coverage while maintaining the smallest prediction set sizes, which highlights the potential of our method to achieve optimal performance by leveraging the strengths of multiple non-conformity scores. Future research directions include exploring the application of our approach to other domains and investigating the integration of advanced machine learning techniques to further enhance the performance and efficiency of conformal prediction sets.
Appendix A Proof of Lemma 1
A.1 VC Dimension of the Subgraph Classes
@@
A.2 Proof of the Lemma
This proof basically follows from Vapnik–Chervonenkis theory. We will use the theorems and lemmas in Wainwright (2019) as reference. The proof of Eq. (9) and Eq. (10) are almost identical. We will focus on Eq. (9). Let $\epsilon_{i},i\in\mathcal{I}$ be i.i.d. symmetric random variables takes value $-1$ or $1$, i.e., $\mathbb{P}(\epsilon_{i}=-1)=\mathbb{P}(\epsilon_{i}=1)=0.5.$ Then we define the Rademacher complexity for the function class in Eq. (11),

### v2

@@ 3.5 Consistency of DLCP
In the setting of DLCP, $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We will denote this by $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ respectively. In this case, the test set and $\widehat{\bm{w}}$ become dependent. The prediction set size is presented differently than previous results.
@@
Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 for a test sample $x$ with the setting of DLCP. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}$ $\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}$
$1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
@@ 3.6 Consistency of DLCP+
Similar as the DLCP setting, $\widehat{\bm{w}}$ depends on the test set, so the result of prediction set size is presented in a similar way. In the following theorem, we use $\mathcal{I}_{2}$ to denote $\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$.
@@
Let $\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x,\widehat{\bm{w}}))$ be the output of Algorithm 2 with the setting of DLCP+. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}$
$\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for test sample $X$,
@@
3.7 Conclusion of Theoretical Results
We integrate the results of Lemma 1 with our main theorems. The Lemma 1 indicates that as the size of the set $\mathcal{I}$ approaches infinity, the events $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with probability $1-o(1)$ even if $\eta,\xi\to 0$ is at a sufficient slow rate. Consequently, the terms $\eta_{i}$ and $\xi_{i}$ in our theorems become negligible. This suggests that the proposed methods achieve coverage rates close to $1-\alpha$ and exhibit near-optimal efficiency.
@@ 4.1 Score Weighting
We conducted an experiment to compare the performance of various score functions score functions discussed in Section 2.2: THR Sadinle et al. (2019), APS Romano et al. (2020), and RANK Luo and Zhou (2024) given by a pretrained classifier $\widehat{p}_{y}(x)$. Additional details about implementation can be found in Section E in the appendix. Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}_{y}(x)$ was available, and the split of the dataset unseen during the training of the pretrained classifier.
@@
Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$ on CIFAR-100. Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size.
@@
In the experiments on CIFAR-10 and CIFAR-100, testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets. The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05. 100 runs with different index splits were conducted to ensure robustness. We have additional experiments on data splitting ratio, which can be found in Section F in the supplimentary file.
@@
Figure 3: Comparison of size vs. coverage for various score functions and our proposed method across $\alpha$ values (0.01-0.05). Our weighted combination method (red) consistently outperforms the other baseline methods by achieving the desired coverage rate with smaller prediction set sizes.
@@
VFCP 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
EFCP 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
DLCP 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
DLCP+ 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
@@
Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset. Results are shown as mean (standard deviation). The first four methods corresponds to the splitting methods in Section 2.5.
The primary objective of our experiments was to evaluate the performance of the proposed weighted score function in comparison with three foundational base score functions: APS, THR, and RANK. Additionally, we compared our method against two competitive baseline score functions, RAPS Angelopoulos et al. (2021) and SAPS Huang et al. (2024).
Figure 2 compares coverage and prediction set sizes across methods at $\alpha=0.01$ on CIFAR-100, while Figure 3 evaluates performance across $\alpha\in[0.01,0.05]$. Using VFCP splits for consistency, our method achieves guaranteed coverage with the smallest prediction sets overall. The advantages are particularly pronounced on CIFAR-10 across all $\alpha$ values and on CIFAR-100 for $\alpha\leq 0.02$.
Notably, THR exhibits strong performance on less challenging tasks (CIFAR-10), where our weighting scheme naturally assigns it dominant weights. In these cases, score averaging provides limited improvement. However, for complex multi-class scenarios (CIFAR-100) or stringent significance levels ($\alpha<0.05$), our method demonstrates clear superiority by optimally combining score functions.
Furthermore, to explore the influence of different data splitting strategies (as discussed in Section 2.4), we conducted additional comparisons of our weighted score function using various split approaches. We also included the performance results of the five individual score functions when utilizing the VFCP split method. Table 1 provides a comprehensive summary of the coverage and size metrics at significance levels of $\alpha=0.01$ and $\alpha=0.05$. These results highlight the effectiveness of the different data splitting strategies. It is important to note that while APS aims to achieve conditional coverage, the other methods, including our approach, do not specifically target conditional coverage. Consequently, it is not surprising that the alternative methods exhibit better efficiency compared to APS.
4.2 Model Weighting
In addition to choosing the score functions in Section 2.2, we further conducted an experiment to compare the performance of various models and their weighted combinations for prediction set construction on CIFAR-10 and CIFAR-100. For CIFAR-10, we used the models ResNet-56, ShuffleNetV2 (1.0x), and VGG16-BN. For CIFAR-100, we used the models VGG16-BN, RepVGG-A2, and MobileNetV2 (1.0x). The performance of the weighted combination method was compared against individual models.
Specifically, for each of the four score functions: THR, APS, RAPS, and SAPS, we generated prediction sets by combining the scores of different models using weights selected by our proposed method. The results are visualized in Figures 4 and 5, which demonstrate the effectiveness of our weighted combination method in terms of weighing scores of different models.
Figure 4: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-10 dataset across $\alpha$ values (0.01–0.05).
Figure 5: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-100 dataset across $\alpha$ values (0.01–0.05).
@@ 5 Related Work
Our work builds upon existing research in conformal prediction, particularly focusing on model averaging and calibration techniques. In the realm of model aggregation, Yang et al. Yang and Kuchibhotla (2024) introduced two selection algorithms designed to minimize the width of prediction intervals by aggregating and selecting from multiple regression estimators.
The technique of VC dimension has been applied in Yang and Kuchibhotla (2024) and Candès et al. (2023). In Section D.2 of Yang and Kuchibhotla (2024), the uniform probability bound is over a finite set. In Candès et al. (2023), the authors considers maximization over real numbers. They have not consider the techiniques subgraph theory like our approach. The methods proposed in Carlsson et al. (2014); Linusson et al. (2017) can aggregate different prediction models to perform conformal prediction, but the methods do not focus on optimizing the efficiency.
A concurrent preprint Liang et al. (2024) examines parametrized score functions $s_{\lambda}(x,y)$ for $\lambda\in\Lambda$ in regression settings. While their general framework handles arbitrary parameter spaces, controlling the Rademacher complexity becomes challenging except in specialized cases. Our approach differs fundamentally by combining pre-defined score functions linearly. This structural constraint enables theoretical guarantees dependent only on the number of constituent scores rather than the complexity of function spaces.
6 Conclusion and Discussion
Our weighted score aggregation method enables efficient and valid prediction set construction for multi-class classification through optimized combinations of score functions and strategic data splitting. Theoretically, we establish finite-sample coverage guarantees and oracle inequalities quantifying the efficiency gap between our method and optimal weights. Empirically, experiments demonstrate consistent maintenance of coverage requirements with minimal prediction set sizes compared to single-score baselines. This work bridges model averaging and conformal prediction, providing a flexible framework for uncertainty quantification that adapts to dataset characteristics through optimal score combinations.
These results suggest several directions for future work:
  1. 1.
    Regression Extension: Algorithmic adaptation to regression requires new theoretical tools due to continuous output spaces, where finite-class union bounds become inapplicable. Potential approaches include metric entropy analysis or covering number techniques.
  2. 2.
    Optimization Enhancement: Developing gradient-based alternatives to grid search, such as differentiable conformal objectives or online weight adaptation during model training, would improve scalability with many score functions.
Appendix A Grid Search Implementation
To solve Eq. (4), we discretize the probability simplex:
$\Delta^{d-1}=\left\{\bm{w}\in\mathbb{R}_{+}^{d}:\sum_{j=1}^{d}w_{j}=1\right\}$
using a grid resolution $\varepsilon=0.01$. Candidate weights are generated as:
$\mathcal{W}=\left\{\bm{w}=(k_{1}\varepsilon,\ldots,k_{d}\varepsilon)\,\,\bigg{|}\,\,k_{j}\in\mathbb{N},\,\,\sum_{j=1}^{d}k_{j}=\lfloor 1/\varepsilon\rfloor\right\},$
yielding $\binom{\lfloor 1/\varepsilon\rfloor+d-1}{d-1}$ distinct weight vectors. For $d=3$ scores, this produces 5,151 candidates, ensuring comprehensive coverage of the parameter space while remaining computationally tractable through parallel evaluation.
Appendix B Proof of Lemma 1
B.1 VC Dimension of the Subgraph Classes
@@
B.2 Proof of the Lemma
This proof basically follows from Vapnik–Chervonenkis theory. We will use the theorems and lemmas in Wainwright (2019) as reference. The proof of both parts of the lemma are almost identical. We will focus on part (a). Let $\epsilon_{i},i\in\mathcal{I}$ be i.i.d. symmetric random variables takes value $-1$ or $1$, i.e., $\mathbb{P}(\epsilon_{i}=-1)=\mathbb{P}(\epsilon_{i}=1)=0.5.$ Then we define the Rademacher complexity for the function class in Eq. (9),